OpenStudy (anonymous):

What is the volume of a cube, if the length of one side is 9 inches?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

nvm take 81(9)

OpenStudy (anonymous):

you get 729

OpenStudy (anonymous):

Volume=lengthxwidthxheight but since its a cube and its all the same length volume of a cube is v=s^3 so v=9^3 v=729in^3

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^ ahh yea the ^3 :)